OpenStudy (anonymous):

1.Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3) A. y equals negative two-thirds x plus 3 B. y equals negative two-thirds x plus five-thirds C. y = 2x – 21 D. y equals negative two-thirds x minus 9 2.Write the equation of a line in slope-intercept form that has a slope of -2 and passes through the point (2, -8). A. y = -2x - 4 B. y = -2x - 8 C. y = -2x + 18 D. y = -2x – 12

OpenStudy (anonymous):

3.Write an equation of a line in point-slope form that has a slope of -2 and passes through (5, -1). A. y + 1 = -2(x – 5) B. y – 1 = -2(x – 5) C. y – 5 = -2(x + 1) D. y -5 = -2(x – 1) Please help, i am pretty much failing everything. I have been very sick and just started this year and need to get caught up bad!

OpenStudy (anonymous):

\[2x+3y=5\] \[3y=5-2x\] \[y=-\frac{ 2 }{ 3 }x-\frac{ 5 }{ 3 }\] So the slope is going to be -3/5. Now plug in your point values for x and y into: \[y=-\frac{ 2 }{ 3 }x+b\] \[-3=-\frac{ 2 }{ 3 }*9+b\] \[-3=-6+b\] \[3=b\] So the y intercept is 3 and the equation is \[y=-\frac{ 2 }{ 3 }x+3\]

OpenStudy (anonymous):

Thank You